# MoralCLIP: Contrastive Alignment of Vision-and-Language Representations with Moral Foundations Theory

### v1

@@ 3.3. Morally-Grounded Data Augmentation
Visual Moral Classifier.    The Visual Moral Compass is a high-quality image moral classifier, built on a fine-tuned CLIP (ViT-B/16) vision encoder, followed by five independent classifier heads—one for each moral foundation. Only the final encoder layer and classifier heads are trained. Each head outputs a probability distribution over three classes: Virtue, Vice, or Neither, enabling multi-label classification that reflects the pluralistic nature of moral judgment (Graham et al., 2013). The classifier is trained on a preprocessed version of SMID, where each image is labeled according to its moral dimension within a given foundation (Appendix A). For loss formulation and implementation details, refer to Appendix B.
@@
Moral Captions.    When generating captions for morally negative images, most models tend to refuse responses or sanitize descriptions (Beyer et al., 2024; Dai et al., 2023). To mitigate this limitation, we used the MoonDream2B captioning model (Vik Korrapati, 2024) [^2] Model available at HuggingFace, version from 26-08-2024. which can generate accurate depictions of morally negative scenes—an essential property for our inference setting. Ten captions per image were generated for the ImageNet and SMID datasets, and five captions per image for the LAION dataset due to its larger size. The captions are concise descriptions of the images, making them well-suited for CLIP’s context window.
@@
Agreement.    This pipeline resulted in a weakly labeled dataset of paired images and captions, each annotated with predicted moral foundation labels. To validate our automated labeling approach, we conducted human evaluation (see Section 5.2 for validation results).
@@
Table 2. Performance comparison of different MoralCLIP variants across modalities on the SMID subset of the test set. We report Mean Average Precision (MAP) to measure retrieval performance of morally similar content, Discriminative Power (DP) to quantify separation between moral categories, and Silhouette Score to assess embedding separation quality. For cross-modal evaluation we use MAP for both image-to-text (I2T) and text-to-image (T2I) retrieval. Best performing model is highlighted in bold. M and S denote Mild and Strong MFT mixing strategies, respectively. All metrics are reported with standard errors computed via bootstrap resampling ($n=1000$).
@@
MoralCLIP Normal ($\lambda=0.5$) $65.51\pm 2.58$ $1.160\pm 0.008$ $0.084\pm 0.014$ $58.61\pm 1.96$ $1.071\pm 0.006$ $0.048\pm 0.011$ $63.88\pm 1.75$ $65.73\pm 2.31$
@@ 4.1. Baselines
Augmented.    Enhances the training with standard data augmentation (e.g., resized crops and random transformations) [^3] Specific augmentation techniques include random horizontal flips, color jittering, and random resized crops.. Each training sample is augmented four times, with each augmented version paired with alternative—but still semantically accurate—captions from our dataset. These augmentations preserve moral content while varying visual properties, testing whether visual robustness alone improves generalization in moral grounding.
MFT Swapper.    Building upon the Augmented variant, this version implements content mixing between samples that share moral dimensions for 75% of the dataset. Specifically, we randomly swap images and text descriptions between samples with matching moral labels, creating new image-text pairs while preserving their moral associations. This approach assesses if the model can learn moral concepts that generalize further than specific image-text instances—for example, whether it can recognize that different manifestations of Care, such as helping elderly, sharing food, caring for animals, represent the same underlying moral dimension. By exposing the model to diverse expressions of the same moral concepts, this variant aims to improve generalization in moral understanding. We explore two levels of this mixing strategy: (1) Mild, which limits the number of swaps per moral dimension group to ensure uniform sampling across moral concepts; and (2) Strong, which allows more frequent moral labels to be sampled more often.
This experimental design allows us to systematically evaluate different approaches to improving moral understanding in multimodal models. Training each variant under both implicit and explicit supervision enables us to assess whether these enhancements are more effective when moral learning is guided explicitly or emerges implicitly through contrastive training. Both visual and text encoder models were initialized from CLIP’s ViT-B/16, with full fine-tuning of all layers during the contrastive moral alignment training. Complete implementation details for all variants are provided in Appendix D.
@@ 4.2. Datasets
The total number of samples in our dataset $\mathcal{D}$ is 15,000 morally labeled image-text pairs. We used standard held-out splits for evaluation: 5% validation and 5% test splits from all three source datasets for MoralCLIP evaluation. Both splits preserve the relative distribution of moral foundations to ensure balanced evaluation.
For model selection, we exclusively used the SMID portion of the validation set for computing metrics, as it contains the strongest moral signal with expert curated annotations. For final evaluation, we used different test set portions depending on analysis. Retrieval analysis used exclusively queries from the SMID test subset to leverage its reliable moral signal, while embedding visualization analyses used the complete test set to provide a comprehensive view across different data distributions.
Figure 3. t-SNE visualization of embedding spaces across different models and modalities. Points are colored by moral categories. Note that the moral annotations are multi-label, meaning individual samples can exhibit multiple moral dimensions simultaneously. Across both image and text embedding spaces, MoralCLIP demonstrates clearer separation between moral categories and better clustering of morally similar content compared to the baseline CLIP model.
@@ 4.3. Methodology and Metrics
To assess the performance of MoralCLIP and other baselines, we developed a multi-criteria evaluation framework that assesses performance across modalities (visual, text, and cross-modal) with three distinct metrics:
Mean Average Precision (MAP).    MAP evaluates retrieval performance by measuring the model’s capability to rank morally similar content based on similarity scores in the embedding space. For each query item, we compute cosine similarities with all other items and rank them in descending order. We consider retrieved items relevant if they share at least one moral label with the query. MAP is particularly suitable for our setting, as it rewards models that retrieve morally aligned content in earlier positions of the similarity-based ranking.
Discriminative Power (DP).    DP quantifies moral category separation by computing the ratio of intra-class similarity (items which share at least one moral label) to inter-class similarity (items with no shared labels). Higher values indicate stronger within-class coherence and greater between-class separation.
Silhouette Score.    Silhouette score assesses whether embeddings are grouped by moral polarity. We simplify our multi-label space to three broader categories (Virtue, Vice, and Neutral) to test whether the embedding space reflects these core moral distinctions.
@@ 5.1. MoralCLIP
Table 2 presents a comprehensive evaluation of MoralCLIP variants across multiple metrics and modalities. These results demonstrate that CLIP + Moral Images outperforms both CLIP and Safe-CLIP ($39.82\%$ versus $41.53\%$ versus $40.51\%$ in terms of I2T MAP), underscoring the importance of a simple implicit moral supervision. This suggests that morally diverse training naturally induces latent moral structure, with stronger effects when we mix captions and images from the same category (MFT Swapper). Notably, Safe-CLIP—a fine-tuned CLIP-Large-14 model—fails to achieve meaningful moral organization despite its larger architecture. This may suggest that Safe-CLIP’s methodology designed to map inappropriate content to safer embedding regions inadvertently diminishes the moral distinctions that our approach seeks to preserve and enhance.
@@
MoralCLIP Augmented, with explicit moral alignment, consistently outperforms implicit approaches and surpasses the safety-focused Safe-CLIP baseline, achieving the best overall performance. This variant shows remarkable gains across all metrics, with MAP scores increasing by 29.62 percentage points for images (from 42.06% to 71.68%) and 23.04 for text (from 38.73% to 61.77%) when compared to the pre-trained version. This improvement indicates that integrating moral similarity directly into the contrastive learning objective effectively reorganizes the embedding space along moral dimensions. This reorganization is visually confirmed through our analysis of the embedding space, which shows that MoralCLIP creates more coherent moral clusters (see Section 5.3.1).
A closer comparison across model types reveals that the optimal strategy differs between training paradigms. In the implicit setting, MFT Swapper variants outperform all baselines, suggesting that mixing content within moral categories helps the model discover latent moral structure. However, with explicit supervision, standard augmentation surpasses MFT Swapper. This indicates that when moral labels are provided explicitly, maintaining semantic coherence matters more than moral diversity. This finding aligns with Park et al. (2024), who show that moral pluralism is challenging to deduce via self-supervision alone and typically requires explicit labels. Our results similarly suggest that explicit moral supervision enables clearer distinction between moral dimensions.
@@
Across all models and metrics, the image modality consistently outperforms text. This advantage may stem from visual content often containing more explicit moral cues than text (Keib et al., 2018; Crone et al., 2018; McCurrie et al., 2018; Bavel and Robertson, 2024) and limitations of CLIP’s text encoder, which include a short effective input length (Zhang et al., 2025) and limited capacity for processing subtleties in language (Li et al., 2025). This asymmetry suggests that visual content may serve as a stronger signal for moral alignment, while simultaneously highlighting the need for improved text encoders capable of handling longer, subtler moral narratives. Despite this modality imbalance, our cross-modal results demonstrate that MoralCLIP achieves robust bidirectional alignment—an essential property for consistent moral interpretation across input modalities.
@@
Figure 4. Image-to-Image retrieval comparison between MoralCLIP and CLIP models on the test set. MoralCLIP retrieves similar images depicting human connection across diverse contexts, while CLIP focuses on low-level visual features like color scheme and formal posing. To ensure consistent evaluation across modalities, we use the same query across all four modality pairs (see Appendix F). Similarity scores represent cosine similarity. The moral labels in bold match the query’s label.
@@ 5.2. Morally-Grounded V&L Data Augmentation
To augment the dataset of image-text pairs with moral labels, we leveraged the Visual Moral Compass classifier (Section 3.3), which demonstrates strong and consistent performance across various moral foundations, enabling scalable and effective automated labeling for MoralCLIP training (see Appendix C for detailed metrics).
To assess the reliability of these automated annotations, we conducted a human annotation study using a subset of 200 images from our dataset. Twelve annotators participated across four batches of 50 images each, with three annotators per batch. One annotator was excluded due to insufficient response variability ($\sigma$=0.032). We report inter-annotator agreement using Krippendorff’s $\alpha$, and human agreement using Cohen’s $\kappa$ against majority labels (Table 3).
Table 3. Inter-annotator agreement (Krippendorff’s $\alpha$) and human-classifier agreement against majority vote (Cohen’s $\kappa_{\text{maj}}$). Consensus Coverage (CC) shows the percentage of examples with annotator consensus.
Foundation IAA ($\alpha$) Model ($\kappa_{\text{maj}}$) CC
@@
The results reveal substantial variation in human agreement across moral foundations ($\alpha=0.184-0.417$), with Care showing the highest consensus and Fairness the lowest—a pattern that closely mirrors our classifier’s performance hierarchy (Appendix C). These trends are consistent with findings in moral psychology that suggest certain foundations are less culturally variable than others (Graham et al., 2013, 2011). For instance, Care is often regarded as a universal moral concern, whereas Fairness and Authority show greater cultural differences and ideological divergence (Atari et al., 2023; Graham et al., 2013, 2011). Moreover, the agreement levels we observe are consistent with—and in some cases exceed—those reported in prior moral annotation efforts, including MFTC (Hoover et al., 2020) and MFRC (Trager et al., 2022).
Importantly, model-human agreement followed a similar pattern, with highest alignment on Care ($\kappa_{\text{maj}}=0.451$) and lowest on Authority ($\kappa_{\text{maj}}=0.024$), reflecting the differing levels of annotator consensus. These results, combined with high consensus coverage (83.7%–94.6%), validate that our classifier captures genuine moral patterns suitable for MoralCLIP training. Additional details on our annotation tool and protocol are available in Appendix E.
@@ 5.3.1. Embedding Space Visualization
In contrast, MoralCLIP-Augmented (second and fourth panels) displays improved moral organization with clear virtue-vice separation across both modalities. This moral clustering effect is consistent across both modalities, though more pronounced in the image embedding space than in text, aligning with our quantitative findings (Table 2). This visualization indicates that moral supervision successfully transforms semantically-organized representations into ones that are organized along moral dimensions.
@@ 5.3.2. Retrieval Analysis
To further investigate how MoralCLIP’s embedding space differs from standard CLIP, we analyze retrieval performance using MoralCLIP-Augmented. Using a consistent query across all four modality combinations, Figure 4 depicts the image-to-image retrieval comparison and Figure 5 depicts text-to-image retrieval. For complete examples refer to Appendix F.
The results reveal a fundamental difference in how the two models interpret content across modalities. We illustrate this with a query depicting a handshake, symbolizing cooperation and respect. In image-to-image retrieval, MoralCLIP retrieves diverse yet morally aligned content: dog companionship (nurturance), collaborative labor (cooperation), two individuals in conversation (connection), and a contemplative military serviceman (respect). Similarly, in text-to-image retrieval (Figure 5), MoralCLIP retrieves images emphasizing moral themes of care and human connection rather than literal visual matches. This demonstrates that MoralCLIP’s similarity metric is driven by moral semantic associations rather than surface-level patterns, with all retrieved content showing significant moral label overlap with the query. In contrast, CLIP’s results clearly reflect its reliance on surface-level characteristics across both tasks. In image retrieval, all results are black-and-white, suggesting the model prioritizes the monochromatic nature of the query over its semantic content. In text-to-image retrieval, CLIP similarly retrieves black-and-white historical imagery, responding to the ”historic” descriptor mentioned in the query text. This aligns with prior findings that CLIP associates grayscale imagery with historical contexts (Offert, 2023; Barancová et al., 2023) and struggles with fine-grained detail (Krojer et al., 2022; Bianchi et al., 2024). While CLIP’s literal interpretation is technically accurate, it focuses on descriptive and visual characteristics rather than the underlying moral significance of human interaction, missing the deeper themes of cooperation and respect that transcend historical context.

### v2

@@ 3.3. Morally-Grounded Data Augmentation
Visual Moral Classifier.   The Visual Moral Compass is a high-quality image moral classifier, built on a fine-tuned CLIP (ViT-B/16) vision encoder, followed by five independent classifier heads—one for each moral foundation (architecture and training details in Appendix B). Only the final encoder layer and classifier heads are trained. Each head outputs a probability distribution over three classes: virtue, vice, or neutral, enabling multi-label classification that reflects the pluralistic nature of moral judgment (Graham et al., 2013). The classifier is trained on a preprocessed version of SMID consisting of 2,401 entries, where each image is labeled according to its moral dimension within a given foundation. Detailed preprocessing steps are described in Appendix A.
@@
Moral Captions.   When generating captions for morally negative images, most models tend to refuse responses or sanitize descriptions (Beyer et al., 2024; Dai et al., 2023). To mitigate this limitation, we used the MoonDream2B captioning model (Vik Korrapati, 2024) [^2] Model available at HuggingFace, version from 26-08-2024. which can generate accurate depictions of morally negative scenes—an essential property for our setting. Ten captions per image were generated for the ImageNet and SMID datasets, and five captions per image for the LAION dataset due to its larger size. The captions are concise descriptions of the images, making them well-suited for CLIP’s context window.
@@
Agreement.   This pipeline resulted in a weakly labeled dataset of paired images and captions, each annotated with predicted moral foundation labels. To validate our automated labeling approach, we conducted a human evaluation (see Section 5.2 for summary results; detailed annotation procedure and interface are provided in Appendix E, and full classifier metrics in Appendix C). .
@@
Table 2. Performance comparison of MoralCLIP variants across modalities on the SMID test subset. We report Mean Average Precision (MAP) to measure retrieval performance of morally similar content, Discriminative Power (DP) to quantify separation between moral categories, and Silhouette Score to assess embedding separation quality. For cross-modal evaluation we use MAP for both image-to-text (I2T) and text-to-image (T2I) retrieval. Best performing models are in bold. M and S denote Mild and Strong MFT mixing strategies, respectively. All metrics include standard errors computed via bootstrap resampling ($n=1000$).
@@
MoralCLIP
Normal ($\lambda=0.5$) $65.51\pm 2.58$ $1.160\pm 0.008$ $0.084\pm 0.014$ $58.61\pm 1.96$ $1.071\pm 0.006$ $0.048\pm 0.011$ $63.88\pm 1.75$ $65.73\pm 2.31$
@@ 4.1. Baselines
Augmented.   Enhances the training with standard data augmentation [^3] Specific augmentation techniques include random horizontal flips, color jittering, and random resized crops.. Each training sample is augmented four times, with each augmented version paired with alternative—but still semantically accurate—captions from our dataset. These augmentations preserve moral content while varying visual properties, testing whether visual robustness alone improves generalization in moral grounding.
MFT Swapper.   Building upon the Augmented variant, this version implements content mixing between samples that share moral dimensions for 75% of the dataset. Specifically, we randomly swap images and text descriptions between samples with matching moral labels, creating new image-text pairs while preserving their moral associations. This approach assesses if the model can learn moral concepts that generalize further than specific image-text instances—for example, whether it can recognize that different manifestations of Care, such as helping the elderly, sharing food, caring for animals, represent the same underlying moral dimension. We explore two levels of this mixing strategy: (1) Mild, which limits the number of swaps to a maximum of 500 per moral dimension group to ensure uniform sampling across moral concepts; and (2) Strong, which allows more frequent moral labels to be sampled more often.
This experimental design allows us to systematically evaluate different approaches to improving moral understanding in multimodal models. Training each variant under both implicit (Section 3.1) and explicit (Section 3.2) supervision enables us to assess whether these enhancements are more effective when moral learning is guided explicitly or emerges implicitly through contrastive training. Both visual and text encoder models were initialized from CLIP’s ViT-B/16, with full fine-tuning of all layers during the contrastive moral alignment training.
@@ 4.2. Datasets
The total number of samples in our dataset $\mathcal{D}$ is 15,000 morally labeled image-text pairs. We used standard held-out splits for evaluation: 5% validation and 5% test splits from all three source datasets for MoralCLIP assessment. Both splits preserve the relative distribution of moral foundations to ensure balanced evaluation.
For model selection, we exclusively used the SMID portion of the validation set for computing metrics, as it contains the strongest moral signal with expert-curated annotations. For final evaluation, we used different test set portions depending on the analysis. Retrieval analysis used exclusively queries from the SMID test subset to leverage its reliable moral signal, while embedding visualization analyses used the complete test set to provide a comprehensive view across different data distributions.
Figure 3. t-SNE visualization of embedding spaces across different models and modalities. Points are colored by moral categories. Note that the moral annotations are multi-label, meaning individual samples can exhibit multiple dimensions simultaneously. Across both image and text embedding spaces, MoralCLIP demonstrates clearer separation between moral categories and better clustering of morally similar content compared to the baseline CLIP model.
@@ 4.3. Methodology and Metrics
To evaluate the performance of MoralCLIP and other baselines, we developed a multi-criteria evaluation framework that assesses performance across modalities with three distinct metrics:
Mean Average Precision (MAP).   MAP (Manning et al., 2008) evaluates retrieval performance by measuring the model’s capability to rank morally similar content based on similarity scores in the embedding space. For each query item, we compute cosine similarities with all other items and rank them in descending order. We consider retrieved items relevant if they share at least one moral label with the query. MAP is particularly suitable for our setting, as it rewards models that retrieve morally aligned content in earlier positions of the similarity-based ranking.
Discriminative Power (DP).   DP quantifies moral category separation by computing the ratio of intra-class similarity (items which share at least one moral label) to inter-class similarity (items with no shared labels). Higher values indicate stronger within-class coherence and greater between-class separation (Wu et al., 2017).
Silhouette Score.   Silhouette score (Rousseeuw, 1987) assesses whether embeddings are grouped by moral polarity. We simplify our multi-label space to three broader categories (virtue, vice, and neutral) to test whether the embedding space reflects these core moral distinctions.
@@ 5.1. MoralCLIP
Table 2 presents a comprehensive evaluation of MoralCLIP variants across multiple metrics and modalities. These results demonstrate that CLIP + Moral Images (the implicit setting, Section 3.1) outperforms both CLIP and Safe-CLIP, underscoring the importance of simple implicit moral supervision. This suggests that morally diverse training naturally induces latent moral structure, with stronger effects when we mix captions and images from the same category (MFT Swapper). Notably, Safe-CLIP—a fine-tuned CLIP-Large-14 model—fails to achieve meaningful moral organization despite its larger architecture. This may suggest that Safe-CLIP’s methodology designed to map inappropriate content to safer embedding regions inadvertently diminishes the moral distinctions that our approach seeks to preserve and enhance.
@@
MoralCLIP Augmented, with explicit moral alignment, consistently outperforms implicit approaches and surpasses the safety-focused Safe-CLIP baseline, achieving the best overall performance. This variant shows remarkable gains across all metrics, with MAP scores increasing by 29.62 percentage points for images (from 42.06% to 71.68%) and 23.04 for text (from 38.73% to 61.77%) when compared to the pretrained version. This improvement indicates that integrating moral similarity directly into the contrastive learning objective effectively reorganizes the embedding space along moral dimensions. This reorganization is visually confirmed through our analysis of the embedding space, which shows that MoralCLIP creates more coherent moral clusters (Figure 3). While silhouette scores remain modest—reflecting the context-dependent nature of moral content—the substantial improvements in MAP and DP metrics demonstrate that MoralCLIP effectively captures moral similarity relationships.
Figure 4. Image-to-Image retrieval comparison between MoralCLIP and CLIP models on the test set. MoralCLIP retrieves similar images depicting human connection across diverse contexts, while CLIP focuses on low-level visual features like color scheme and formal posing. Similarity scores represent cosine similarity. The moral labels in bold match the query’s label. Throughout our figures, values marked with * indicate rounding approximations where actual values differ slightly.
@@
A closer comparison across model types reveals that the optimal strategy differs between training paradigms. In the implicit setting (CLIP + Moral Images), MFT Swapper variants outperform all baselines, suggesting that mixing content within moral categories helps the model discover latent moral structure. However, with explicit supervision (MoralCLIP), standard augmentation surpasses MFT Swapper. This indicates that when moral labels are provided explicitly, maintaining semantic coherence matters more than moral diversity. This finding aligns with Park et al. (2024), who show that moral pluralism is challenging to deduce via self-supervision alone and typically requires explicit labels. Our results similarly suggest that explicit moral supervision enables clearer distinction between moral dimensions. Across all models and metrics, the image modality consistently outperforms text. This pattern aligns with the well-documented ”modality gap”, a spatial separation between image and text embeddings that emerges inherently from contrastive training (Liang et al., 2022; Eslami and de Melo, 2025), suggesting that our moral alignment training, while beneficial overall, may have amplified this modality imbalance. This may also stem from visual content often containing more explicit moral cues than text (Keib et al., 2018; Crone et al., 2018; McCurrie et al., 2018; Bavel and Robertson, 2024) and limitations of CLIP’s text encoder, which include a short effective input length (Zhang et al., 2025) and limited capacity for processing subtleties in language (Li et al., 2025). This asymmetry suggests that visual content may serve as a stronger signal for moral alignment, while simultaneously highlighting the need for improved text encoders capable of handling longer, subtler moral narratives. Despite this modality imbalance, our cross-modal results demonstrate that MoralCLIP achieves robust bidirectional alignment—an essential property for consistent moral interpretation across input modalities.
@@ 5.2. Morally-Grounded V&L Data Augmentation
To augment the dataset of image-text pairs with moral labels, we leveraged the Visual Moral Compass classifier (Section 3.3), which demonstrates strong and consistent performance across various moral foundations, enabling scalable and effective automated labeling for MoralCLIP training.
To assess the reliability of these automated annotations, we conducted a human annotation study using a subset of 200 images from our dataset. Twelve annotators participated across four batches of 50 images each, with three annotators per batch. One annotator was excluded due to insufficient response variability ($\sigma$=0.179). We report inter-annotator agreement using Krippendorff’s $\alpha$, and human-classifier agreement using Cohen’s $\kappa$ against majority labels (Table 3).
The results reveal substantial variation in human agreement across moral foundations ($\alpha=0.184-0.417$), with Care showing the highest consensus and Fairness the lowest—a pattern that closely mirrors our classifier’s performance hierarchy, where Care achieved the strongest F1 ($0.84$) and Fairness the weakest ($0.71$), with other foundations clustering around $0.81$. These trends are consistent with findings in moral psychology that suggest certain foundations are less culturally variable than others (Graham et al., 2013, 2011). For instance, Care is often regarded as a universal moral concern, whereas Fairness and Authority show greater cultural differences and ideological divergence (Atari et al., 2023; Graham et al., 2013, 2011). Moreover, the agreement levels we observe are consistent with—and in some cases exceed—those reported in prior moral annotation efforts, including MFTC (Hoover et al., 2020) and MFRC (Trager et al., 2022).
Table 3. Inter-annotator agreement (Krippendorff’s $\alpha$) and human-classifier agreement against majority vote (Cohen’s $\kappa_{\text{maj}}$). Consensus Coverage ($C$) shows the percentage of examples with annotator consensus.
Foundation IAA ($\alpha$) Model ($\kappa_{\text{maj}}$) $C$
@@
More importantly, model-human agreement followed a similar pattern, with highest alignment on Care ($\kappa_{\text{maj}}=0.451$) and lowest on Authority ($\kappa_{\text{maj}}=0.024$), reflecting the differing levels of annotator consensus. These results, combined with high consensus coverage (83.7%–94.6%), validate that our classifier captures genuine moral patterns suitable for MoralCLIP training. Detailed annotation workflow and instructions are available in Appendix E.
@@ 5.3.1. Embedding Space Visualization
In contrast, MoralCLIP-Augmented (second and fourth panels) displays improved moral organization with clear virtue-vice separation across both modalities. Although consistent, this moral clustering effect is more pronounced in the image embedding space than in text, aligning with our quantitative findings (Table 2). This visualization indicates that moral supervision successfully transforms semantically-organized representations into ones that are organized along moral dimensions.
@@ 5.3.2. Retrieval Analysis
To further investigate how MoralCLIP’s embedding space differs from standard CLIP, we analyze retrieval performance using MoralCLIP-Augmented. Using a consistent query across all four modality combinations, Figure 4 depicts the image-to-image retrieval comparison and Figure 5 showcases the text-to-image retrieval example. Retrieval is performed using cosine similarity in the learned embedding space without any explicit use of moral labels during the retrieval process. Additional cross-modal retrieval examples are included in Appendix F.
The results reveal a fundamental difference in how the two models interpret content across modalities. We illustrate this with a query depicting a handshake, symbolizing cooperation and respect. In image-to-image retrieval, MoralCLIP retrieves diverse yet morally aligned content: dog companionship (nurturance), collaborative labor (cooperation), two individuals in conversation (connection), and a contemplative military serviceman (respect). Similarly, in text-to-image retrieval (Figure 5), MoralCLIP retrieves images emphasizing themes of care and human connection rather than literal visual matches. This demonstrates that MoralCLIP’s similarity metric is driven by moral associations rather than surface-level patterns, with all retrieved content showing significant moral label overlap with the query. In contrast, CLIP’s results clearly reflect its reliance on surface-level characteristics across both tasks. In image retrieval, all results are black-and-white, suggesting the model prioritizes the monochromatic nature of the query over its semantic content. In text-to-image retrieval, CLIP similarly retrieves black-and-white historical imagery, responding to the ”historic” descriptor mentioned in the query text. This aligns with prior findings that CLIP associates grayscale imagery with historical contexts (Offert, 2023; Barancová et al., 2023) and struggles with fine-grained detail (Krojer et al., 2022; Bianchi et al., 2024). While CLIP’s literal interpretation is technically accurate, it focuses on descriptive and visual characteristics rather than the underlying moral significance of human interaction, missing the deeper themes of cooperation and respect that transcend historical context.